OpenStudy (anonymous):

write in standard form- 5/(3-15i)

OpenStudy (anonymous):

is it \(\large \frac{-5}{3-15i} \) ???

OpenStudy (anonymous):

either way.... just multiply by the conjugate of the denominator to both top and bottom then simplify...

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

what am i doing wrong?

jimthompson5910 (jim_thompson5910):

Use the difference of squares rule \[\Large (x-y)(x+y) =x^2-y^2\] to say that \[\Large (3-15i)(3+15i) =3^2 - (15i)^2\] \[\Large (3-15i)(3+15i) =9 - 225i^2\] \[\Large (3-15i)(3+15i) =9 - 225(-1)\] \[\Large (3-15i)(3+15i) =9 + 225\] \[\Large (3-15i)(3+15i) =234\]

OpenStudy (anonymous):

that makes sense but how do i get to one of these answers?

jimthompson5910 (jim_thompson5910):

The numerator becomes \[\Large -5(3+15i) = -15 - 75i\] So \[\Large \frac{-5}{3-15i}\] turns into \[\Large \frac{-15 - 75i}{234}\] Then what?

OpenStudy (anonymous):

Oh! It's just simplifying from there! Thank you!! :)

jimthompson5910 (jim_thompson5910):

yes pretty much, tell me what you get

OpenStudy (anonymous):

-5/78 - 25/78 i

jimthompson5910 (jim_thompson5910):

Good \[\Large \frac{-5}{3-15i} = -\frac{5}{78} - \frac{25}{78}i\]

OpenStudy (anonymous):

Thanks a bunch! :)

jimthompson5910 (jim_thompson5910):

sure thing